# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v1

@@ Abstract
Evaluating autonomous vehicle stacks (AVs) in simulation typically involves replaying driving logs from real-world recorded traffic. However, agents replayed from offline data do not react to the actions of the AV, and their behaviour cannot be easily controlled to simulate counterfactual scenarios. Existing approaches have attempted to address these shortcomings by proposing methods that rely on heuristics or learned generative models of real-world data but these approaches either lack realism or necessitate costly iterative sampling procedures to control the generated behaviours. In this work, we take an alternative approach and propose CtRL-Sim, a method that leverages return-conditioned offline reinforcement learning within a physics-enhanced Nocturne simulator to efficiently generate reactive and controllable traffic agents. Specifically, we process real-world driving data through the Nocturne simulator to generate a diverse offline reinforcement learning dataset, annotated with various reward terms. With this dataset, we train a return-conditioned multi-agent behaviour model that allows for fine-grained manipulation of agent behaviours by modifying the desired returns for the various reward components. This capability enables the generation of a wide range of driving behaviours beyond the scope of the initial dataset, including those representing adversarial behaviours. We demonstrate that CtRL-Sim can efficiently generate diverse and realistic safety-critical scenarios while providing fine-grained control over agent behaviours. Further, we show that fine-tuning our model on simulated safety-critical scenarios generated by our model enhances this controllability.
@@ 1 Introduction
Recent advances in autonomous vehicles (AVs) have significantly enhanced their ability to safely navigate the complexities of urban driving (Kusano et al., 2023). Despite this progress, ensuring operational safety in complex long-tail scenarios, such as unexpected pedestrian behaviours and distracted driving, remains a significant barrier to widespread AV adoption. Simulation has emerged as a promising tool for efficiently validating the safety of autonomous vehicles (AVs) in these long-tail scenarios. However, a core challenge in developing a simulator for AVs is the need for other agents within the simulation to exhibit realistic and diverse behaviours that are reactive to the AV, while being easily controllable. The traditional approach for evaluating AVs in simulation involves fixing the behaviour of other agents to the behaviours exhibited in pre-recorded driving data. However, this testing approach does not allow the other agents to react to the AV’s actions, which yields unrealistic interactions between the AV and the other agents.
@@
To address the issues inherent in non-reactive log-replay testing, prior work has proposed rule-based methods (Treiber et al., 2000; Kesting et al., 2007) to enable reactive agents. However, the behaviour of these rule-based agents are often unrealistic as the rules impose rigid constraints that fail to capture the nuances and complexities of human driving behaviour. More recently, generative models learned from real-world data have been proposed to enhance the realism of simulated agent behaviours (Suo et al., 2021; Xu et al., 2023b; Igl et al., 2022; Suo et al., 2023; Zhong et al., 2023b; Philion et al., 2023). While these methods improve over rule-based approaches in terms of realism, they are either not easily controllable (Suo et al., 2021; Xu et al., 2023b; Philion et al., 2023) or require costly iterative sampling procedures to control the agent behaviours (Rempe et al., 2022; Zhong et al., 2023b; Suo et al., 2023; Zhong et al., 2023a; Chang et al., 2024). Moreover, these learned methods often assume simplistic agent kinematics models with limited realism.
@@
Figure 1: CtRL-Sim allows controllable agent behavior from existing datasets. This allows users to create more interesting edge cases for testing and evaluating AV planners.
@@
In this paper, we propose CtRL-Sim to address these limitations of prior work. The CtRL-Sim framework utilizes return-conditioned offline reinforcement learning (RL) to enable reactive, closed-loop, controllable, and probabilistic behaviour simulation within a physics-enhanced Nocturne (Vinitsky et al., 2022) environment. We process scenes from the Waymo Open Motion Dataset (Ettinger et al., 2021) through Nocturne to curate an offline RL dataset for training that is annotated with reward terms such as “vehicle-vehicle collision” and “goal achieved”. We propose a return-conditioned multi-agent decision Transformer architecture (Chen et al., 2021) within the CtRL-Sim framework to imitate the driving behaviours in the curated dataset. We then leverage exponential tilting of the predicted return distribution (Lee et al., 2022) as a simple yet effective mechanism to control the simulated agent behaviours. While Lee et al. (2022) exponentially tilts towards more optimal outcomes for the task of reward-maximizing control, we instead propose to exponentially tilt in either direction to provide flexible control over both good and bad simulated driving behaviours.
@@
We show examples of how CtRL-Sim can be used to generate counterfactual scenes when exponentially tilting the different reward axes in Figure 1. For controllable generation, CtRL-Sim simply requires specifying a tilting coefficient along each reward axis, which circumvents the need for costly iterative sampling that is required by prior methods. This enables the generation of a wide range of interesting and realistic driving behaviours beyond the scope of the initial dataset, including targeted adversarial behaviours. CtRL-Sim scenarios are efficiently simulated within our physics-extended Nocturne environment, enabling realistic agent dynamics and physical collisions. We summarize our main contributions: 1. We propose CtRL-Sim, which is to the best of our knowledge, the first framework applying offline RL for controllable and reactive behaviour simulation. Specifically, CtRL-Sim employs exponential tilting of factorized reward-to-go to control different axes of agent behaviours. 2. CtRL-Sim’s architecture is based on an autoregressive discrete, multi-agent Decision Transformer tailored for controllable behaviour simulation. 3. We extend the Nocturne simulator (Vinitsky et al., 2022) with a Box2D physics engine, which facilitates realistic vehicle dynamics and collision interactions.
@@
Through extensive evaluation, we demonstrate the effectiveness of CtRL-Sim at producing controllable and realistic agent behaviours compared to prior methods. We also show that finetuning our model in the Nocturne simulator with simulated CtRL-Sim scenarios enhances its controllability. CtRL-Sim has the potential to serve as a useful framework for enhancing the safety and robustness of AV planner policies through simulation-based training and evaluation.
@@
Figure 2: 2(a) (left) shows an outline of our method. Agent and map data at timestep 0 are encoded and fed through a Transformer encoder as context for the decoder, similar to Philion et al. (2023). Trajectories are arranged first by agents, then by timesteps, embedded, and fed through the decoder. For each agent, we encode $(s_{t},G_{t},a_{t})$ (i.e. state, return-to-go, action) and we predict from these $(G_{t},a_{t},s_{t+1},\dots,s_{T})$ ($t$ is current timestep, $T$ is the max timestep). 2(b) (right) shows the forward pass at inference time. The state predicts the return-to-go. The return-to-go is tilted (i.e. reweighed to encourage specific behaviors) and predicts the action, and the action predicts the next states.
2 Background
@@
Prior methods differ in how the policy is modelled during training and how inference is performed; we refer the reader to Table 3 in the Appendix A for a breakdown of the different approaches. As CtRL-Sim employs a return-conditioned policy for controllable simulation, we briefly present a class of related methods that learn return-conditioned policies (RCPs) (Chen et al., 2021; Lee et al., 2022; Srivastava et al., 2019; Kumar et al., 2019; Emmons et al., 2021; Piché et al., 2022; Meng et al., 2021) for offline RL. RCPs are concerned with learning the joint distribution of actions and returns, such that action sampling is conditioned on the return distribution. Most methods will learn the joint distribution by decomposing it and minimizing the following objective:
$\mathcal{L}(\theta)=\mathbb{E}_{\{s_{t},a_{t},G_{t}\}}\big{[}-\log p_{\theta}(a_{t}|s_{t},G_{t})p_{\theta}(G_{t}|s_{t})\big{]}.$ (1)
Instead of modelling the return distribution, the Decision Transformer (Chen et al., 2021) conditions the learned policy on the maximum observed return, $G_{\text{max}}$ in the dataset. That is, at inference time, the actions are sampled from $a_{t}\sim p_{\theta}(a_{t}|s_{t},G_{\text{max}})$. Lee et al. (2022) instead propose to employ the learned return distribution, combined with exponential tilting, in order to sample high-return actions while remaining close to the empirical distribution. This effectively amounts to sampling actions, where
$a_{t}\sim p_{\theta}(a_{t}|o_{t},G_{t})p_{\theta}(G_{t}|o_{t})\exp(\kappa G_{t}-\eta(\kappa))),$ (2)
where $\kappa$ controls how strongly we want to tilt the distribution towards high returns, and $\eta(.)$ is a normalizing constant, which is typically ignored. CtRL-Sim adopts exponential tilting of the predicted return distribution for each agent to finely control agent behaviour. We also explicitly model future sequences of states.
3 CtRL-Sim
A behaviour simulation model is tasked with autoregressively generating the future rollout of all agents in a given scene. This would allow for reactive closed-loop evaluation of the AV planning stack in simulation. Prior data-driven behaviour simulation models learn to imitate driving behaviours found in trajectory datasets. In this section, we present the proposed CtRL-Sim framework for behaviour simulation. We first introduce CtRL-Sim framework in the single-agent setting, and later on show how it extends to the multi-agent setting. Given the state of an agent $s_{t}$ at timestep $t$ and additional context (e.g., the road structure, the agent’s goal), the behaviour simulation model employs a driving policy $\pi(a_{t}|s_{t},m,s_{G})$ and a forward transition model $\mathcal{P}(s_{t+1}|s_{t},a_{t})$ to control the agent in the scene. Note that $a_{t}$ is the action, $m$ is the map context, and $s_{G}$ is the prescribed goal state for the agent in the scene. Using the simulator defined in Section 3.3, we have access to a physically-realistic forward transition model $\mathcal{P}$. In this work, we are interested in how to model the policy $\pi(a_{t}|s_{t},m,s_{G})$ such that we can imitate the real distribution of driving behaviour, and critically, we can control the agent’s behavior to generate long-tail counterfactual scenes.
3.1 Our Approach to Controllable Simulation via Offline RL
@@
The primary insight of this work is the observation that offline RL can be an effective way to perform controllable simulation. That is, the policy distribution over actions can be tilted at inference time towards desirable or undesirable behaviors by specifying different values of return-to-go $G_{t}$. This requires a different formulation of the policy such that it is conditioned on the return $\pi(a_{t}|s_{t},G_{t},s_{G})$ [^1] Note that we omit the additional context $m$ for brevity.. In Table 3, we outline how different approaches in offline RL have learned return-conditioned policies. In this work, we adopt an approach that learns the joint distribution of returns and actions of an agent in a given dataset. Specifically,
$p_{\theta}(a_{t},G_{t}|s_{t},s_{G})=\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})p_{\theta}(G_{t}|s_{t},s_{G}).$ (3)
We note that Gontier et al. (2023) found it helpful to also utilize a model-based return-conditioned policy, whereby the future state is modelled as part of the joint distribution being learned. This is shown to provide a useful regularizing signal for the policy, even though the future state prediction is not directly used at inference time. In this work, we also found it helpful to employ a model-based return-conditioned policy, where we regularize the learned policy by predicting the full sequence of future states. The final distribution we are aiming to model is thus given by:
$p_{\theta}(s_{t+1:T},a_{t},G_{t}|s_{t},s_{G})=p_{\theta}(s_{t+1:T}|s_{t},s_{G},G_{t},a_{t})\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})p_{\theta}(G_{t}|s_{t},s_{G}).$ (4)
@@
This decomposition provides a principled way to sample actions according to the learned action distribution at inference time. Specifically, we obtain actions by first sampling returns $G_{t}\sim p_{\theta}(G_{t}|s_{t},s_{G})$ and then sampling actions $a_{t}\sim\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})$. This sampling procedure corresponds to the imitative policy since the sampled returns are obtained from the learned density that models the data distribution. Following prior work in offline RL (Lee et al., 2022; Gontier et al., 2023; Piché et al., 2022), we can also sample actions from an exponentially-tilted policy distribution. This is done by sampling the returns from the tilted distribution $G^{\prime}_{t}\sim p_{\theta}(G_{t}|s_{t},s_{G})\exp(\kappa G_{t})$, with $G^{\prime}_{t}$ being the tilted return-to-go and where $\kappa$ represents the inverse temperature and modulates how concentrated the probability density is around certain outcomes; higher values of $\kappa$ concentrate more density around the best outcomes or higher returns, while negative values of $\kappa$ concentrate more density on less favourable outcomes or lower returns.
@@
We are interested in modelling and controlling the different components of the reward function rather than maximizing their weighted sum. For example, in this work, we would like to model an agent’s ability to reach its goal, avoid driving off the road, and avoid colliding with other road users. In general, given $C$ reward components, our objective is to learn policies that are conditioned on all its factored dimensions as this would grant us control over each one at test time. This entails modelling separate return components as $G^{c}_{t}\sim p_{\theta}(G^{c}_{t}|s_{t},s_{G})$, where each component $G^{c}_{t}$ is a numerical return-to-go value. Taking into account this factorization, we reformulate the learned policy to explicitly account for the conditioning on all return components $\pi_{\theta}(a_{t}|s_{t},s_{G},G^{1}_{t},\ldots G^{C}_{t})$. At test time, each return component will be accompanied by its own inverse temperature $\kappa^{c}$ to enable control over each return component. With this new policy formulation, we can sample actions that adhere to different behaviours specified by $\{\kappa^{1},\ldots,\kappa^{C}\}$, as shown in Algorithm 2 in Appendix B.
@@
To implement our framework for agent behaviour in the simulator we further extend the approach presented above for single-agent controllable behaviour simulation to the multi-agent setting. Across all agents we have sets for the joint states $\mathbb{S}_{t}$, future states $\mathbb{S}_{t+1:T}$, goal states $\mathbb{S}_{G}$, actions $\mathbb{A}_{t}$, and returns-to-go $\mathbb{G}_{t}$. The final multi-agent joint distribution we model is:
@@ 3.2 Multi-Agent Behaviour Simulation Architecture
In this section, we introduce the proposed architecture for multi-agent behaviour simulation within the CtRL-Sim framework that is used to parameterize the multi-agent joint distribution presented in Equation (5). Following recent work applying autoregressive Transformers for motion prediction and simulation (Seff et al., 2023; Philion et al., 2023), we propose an encoder-decoder Transformer architecture (Vaswani et al., 2017), as illustrated in Figure 2, where the encoder encodes the initial scene and the decoder autoregressively generates the trajectory rollout for all agents in the scene.
@@
Encoder To encode the initial scene, we first process two input modalities: the initial agent states and goals $(\mathbf{s}_{0},\mathbf{s}_{G})$ and the map context $m$, where $\mathbf{s}_{0}$ is the joint initial state of all agents and $\mathbf{s}_{G}$ is the joint goal state of all agents. Each agent $i$’s initial state information $s^{i}_{0}$, which includes the position, velocity, heading, and agent type, is encoded with a two-layer multi-layer perceptron (MLP). Similarly, each agent’s goal $s^{i}_{G}$, which is represented as the final position, velocity, and heading from the driving trajectory log, is also encoded with a 2-layer MLP. We then concatenate the initial state and goal embedding of each agent and feed this through a linear layer to get per-agent embeddings of size $d$. We additionally apply an additive learnable positional encoding to encode the agents’ identities across the sequence of agent embeddings. The map context is represented as a set of road segments $m:=\{\mathbf{r}_{l}\}_{l=1}^{L}$, where each road segment is defined by a sequence of points $\mathbf{r}_{l}:=(p_{l}^{1},\dots p_{l}^{P})$, where $L$ is the number of road segments and $P$ is the number of points per road segment. We apply a per-point 2-layer MLP to the points of each road segment $\mathbf{r}_{l}$. To produce road segment-level embeddings, we then apply attention-based pooling (Lee et al., 2019) on the embeddings of the points within each road segment, yielding $L$ road segment embeddings of size $d$. The initial agent embeddings and road segment embeddings are then concatenated into a sequence of length $N+M$ and processed by a sequence of $E$ Transformer encoder blocks.
@@
Decoder The proposed decoder architecture models the joint distribution in Equation (5) as a sequence modelling problem, where we model the probability of the next token in the sequence conditioned on all previous tokens $p_{\theta}(x_{t}|x_{<t})$. In this work, we consider trajectory sequences of the form: $x=\langle\dots,(s^{1}_{t},s^{1}_{G}),(G^{1,1}_{t},\dots,G^{C,1}_{t}),a^{1}_{t},...,(s^{N}_{t},s^{N}_{G}),(G^{1,N}_{t},\dots,G^{C,N}_{t}),a^{N}_{t},\dots\rangle$, where $s^{i}_{t}$, $s^{i}_{G}$, and $a^{i}_{t}$ correspond to the state, goal, and action of agent $i$ at timestep $t$, and $G^{c,i}_{t}$ corresponds to the $c$’th factored return component for agent $i$ at timestep $t$. These sequences can be interpreted as an extension of the sequences considered in the Multi-Game Decision Transformer (Lee et al., 2022) to the multi-agent goal-conditioned setting with factorized returns. We emphasize that unlike Decision Transformer (Chen et al., 2021), our model explicitly predicts the return distribution and samples from it at inference time, which enables flexible control over the agent behaviours and circumvents the need to specify an expert return-to-go. We obtain state-goal tuple $(s^{i}_{t},s^{i}_{G})$ embeddings in the same way that $(\mathbf{s}_{0},\mathbf{s}_{G})$ are processed in the encoder. Following recent work that tokenizes driving trajectories (Seff et al., 2023; Philion et al., 2023), we discretize the actions and return-to-gos. For the actions, we discretize the acceleration and steering into 20 and 50 uniformly quantized bins, respectively, yielding 1000 action tokens. For the return-to-gos, we discretize each return-to-go component $G^{c,i}_{t}$ into 350 uniformly quantized bins. We then embed the action and return-to-go tokens with a linear embedding. To each input token, we additionally add two learnable positional embeddings representing the agent identity and timestep, respectively. The tokenized sequence is then processed by a stack of $D$ Transformer decoder layers with a temporally causal mask. To ensure that the model is permutation equivariant to the agent ordering (Ngiam et al., 2022; Girgis et al., 2021), we modify the standard temporally causal mask by additionally enforcing that each agent can only attend to its own action and return-to-go tokens at the present timestep while allowing access to all agents’ state tokens at the present timestep and all agents’ tokens in the past timesteps.
@@
Training Given a dataset of offline trajectories (Section 4.1), we train our model by sampling sequences of length $H\times N\times 3$, where $H$ is a hyperparameter specifying the number of timesteps in the context. The state, return-to-go, and action token embeddings outputs by the decoder are used to predict the next return token, action token, and future state sequence, respectively. We train the return-to-go and action headers with the standard cross-entropy loss function and the future state sequence header with an L2 regression loss function. The final loss function is of the form: $\mathcal{L}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{return-to-go}}+\alpha\mathcal{L}_{\text{state}}$, where $\mathcal{L}_{\text{action}}$ and $\mathcal{L}_{\text{return-to-go}}$ are cross-entropy loss functions, $\alpha$ is a controls the strength of the auxiliary state prediction loss, and $\mathcal{L}_{\text{state}}$ is an L2 loss function.
3.3 Physics-based Simulator
CtRL-Sim extends the Nocturne simulation environment (Vinitsky et al., 2022). Nocturne is a lightweight 2D driving simulator that is built on real-world driving trajectory data from the Waymo Open Motion Dataset (Ettinger et al., 2021). Refer to Appendix C for more detail on how scenes are represented in the Nocturne simulator. We extend Nocturne by integrating a physics engine based on the Box2D library for enabling realistic vehicle dynamics and vehicle collisions. We model the vehicle’s dynamics using basic physics principles, where forces applied to the vehicle are translated into acceleration, influencing its speed and direction, and with frictional forces applied to simulate realistic sliding and adherence behaviors. This extension additionally ensures that an agent’s acceleration, braking, and turn radius are bound by plausible limits and that vehicles can physically collide with each other. Such improvements open the possibility of more accurately simulating complex conditions, such as emergency braking maneuvers, slippery roads, and multi-vehicle collisions.

### v2

@@ Abstract
Evaluating autonomous vehicle stacks (AVs) in simulation typically involves replaying driving logs from real-world recorded traffic. However, agents replayed from offline data are not reactive and hard to intuitively control. Existing approaches address these challenges by proposing methods that rely on heuristics or generative models of real-world data but these approaches either lack realism or necessitate costly iterative sampling procedures to control the generated behaviours. In this work, we take an alternative approach and propose CtRL-Sim, a method that leverages return-conditioned offline reinforcement learning to efficiently generate reactive and controllable traffic agents. Specifically, we process real-world driving data through a physics-enhanced Nocturne simulator to generate a diverse offline reinforcement learning dataset, annotated with various reward terms. With this dataset, we train a return-conditioned multi-agent behaviour model that allows for fine-grained manipulation of agent behaviours by modifying the desired returns for the various reward components. This capability enables the generation of a wide range of driving behaviours beyond the scope of the initial dataset, including adversarial behaviours. We demonstrate that CtRL-Sim can generate diverse and realistic safety-critical scenarios while providing fine-grained control over agent behaviours.
Keywords: Autonomous Driving, Simulation, Offline Reinforcement Learning
@@ 1 Introduction
Recent advances in autonomous driving has enhanced their ability to safely navigate the complexities of urban driving [1]. Despite this progress, ensuring operational safety in long-tail scenarios, such as unexpected pedestrian behaviours and distracted driving, remains a significant barrier to widespread adoption. Simulation has emerged as a promising tool for efficiently validating the safety of autonomous vehicles (AVs) in these long-tail scenarios. However, a core challenge in developing a simulator for AVs is the need for other agents within the simulation to exhibit realistic and diverse behaviours that are reactive to the AV, while being easily controllable. The traditional approach for evaluating AVs in simulation involves fixing the behaviour of agents to the behaviours exhibited in pre-recorded driving data. However, this testing approach does not allow the other agents to react to the AV, which yields unrealistic interactions between the AV and the other agents.
@@
To address the issues inherent in non-reactive log-replay testing, prior work has proposed rule-based methods [2, 3] to enable reactive agents. However, the behaviour of these rule-based agents often lacks diversity and is unrealistic. More recently, generative models learned from real-world data have been proposed to enhance the realism of simulated agent behaviours [4, 5, 6, 7, 8, 9]. While these methods produce more realistic behaviours, they are either not easily controllable [4, 5, 9] or require costly sampling procedures to control the agent behaviours [10, 8, 7, 11, 12].
@@
Figure 1: CtRL-Sim allows for controllable agent behaviour from existing datasets. This allows users to create interesting edge cases for testing and evaluating AV planners.
@@
In this paper, we propose CtRL-Sim to address these limitations of prior work. The CtRL-Sim framework utilizes return-conditioned offline reinforcement learning (RL) to enable reactive, closed-loop, controllable, and probabilistic behaviour simulation within a physics-enhanced Nocturne [13] environment. We process scenes from the Waymo Open Motion Dataset [14] through Nocturne to curate an offline RL dataset for training that is annotated with reward terms such as “vehicle-vehicle collision” and “goal achieved”. We propose a return-conditioned multi-agent autoregressive Transformer architecture [15] within the CtRL-Sim framework to imitate the driving behaviours in the curated dataset. We then leverage exponential tilting of the predicted return distribution [16] as a simple yet effective mechanism to control the simulated agent behaviours. While [16] exponentially tilts towards more optimal outcomes for the task of reward-maximizing control, we instead propose to tilt in either direction to provide control over both good and bad simulated driving behaviours.
@@
We show examples of how CtRL-Sim can be used to generate counterfactual scenes when exponentially tilting the different reward axes in Figure 1. For controllable generation, CtRL-Sim simply requires specifying a tilting coefficient along each reward axis, which circumvents the costly iterative sampling required by prior methods. CtRL-Sim scenarios are simulated within our physics-extended Nocturne environment. We summarize our main contributions: 1. We propose CtRL-Sim, which is, to the best of our knowledge, the first framework applying return-conditioned offline RL for controllable and reactive behaviour simulation. Specifically, CtRL-Sim employs exponential tilting of factorized reward-to-go to control different axes of agent behaviours. 2. We propose an autoregressive multi-agent encoder-decoder Transformer architecture within the CtRL-Sim framework that is tailored for controllable behaviour simulation. 3. We extend the Nocturne simulator [13] with a Box2D physics engine, which facilitates realistic vehicle dynamics and collision interactions.
@@
We demonstrate the effectiveness of CtRL-Sim at producing controllable and realistic agent behaviours compared to prior methods. We also show that finetuning our model in Nocturne with simulated adversarial scenarios enhances control over adversarial behaviours. CtRL-Sim has the potential to serve as a useful framework for enhancing the safety and robustness of AV planner policies through simulation-based training and evaluation.
2 CtRL-Sim
@@
Figure 2: 2(a) (left) The agent and map data at $t=0$ are encoded and fed through a Transformer encoder as context for the decoder, similar to [9]. Trajectories are arranged first by agents, then by timesteps, embedded, and fed through the decoder. For each agent, we encode $(s_{t},G_{t},a_{t})$ (i.e. state, return-to-go, action) and we predict from these $(G_{t},a_{t},s_{t+1},\dots,s_{T})$. 2(b) (right) At inference time, the state predicts the return-to-go. The return-to-go is tilted (i.e., reweighed to encourage specific behaviors) and is used to predict the action, which in turn is used to predict the next states.
In this section, we present the proposed CtRL-Sim framework for behaviour simulation. We first introduce CtRL-Sim in the single-agent setting, and subsequently show how it extends to the multi-agent setting. Given the state of an agent $s_{t}$ at timestep $t$ and additional context (e.g., the road structure, the agent’s goal), the behaviour simulation model employs a driving policy $\pi(a_{t}|s_{t},m,s_{G})$ and a forward transition model $\mathcal{P}(s_{t+1}|s_{t},a_{t})$ to control the agent in the scene. Note that $a_{t}$ is the action, $m$ is the map context, and $s_{G}$ is the prescribed goal state. Using the physics-extended Nocturne simulator, we have access to a physically-realistic forward transition model $\mathcal{P}$. In this work, we are interested in modelling the policy $\pi(a_{t}|s_{t},m,s_{G})$ such that we can both imitate the real distribution of driving behaviour and control the agent’s behavior to generate long-tail counterfactual scenes.
2.1 Our Approach to Controllable Simulation via Offline RL
@@
The primary insight of this work is the observation that offline RL can be an effective way to perform controllable simulation. That is, the policy distribution over actions can be tilted at inference time towards desirable or undesirable behaviors by specifying different values of return-to-go $G_{t}$. This requires a different formulation of the policy such that it is conditioned on the return $\pi(a_{t}|s_{t},G_{t},s_{G})$ [^1] Note that we omit the additional context $m$ for brevity.. In Table 3, we outline how different approaches in offline RL have learned return-conditioned policies. In this work, we adopt an approach that learns the joint distribution of returns and actions of an agent in a given dataset. Specifically, $p_{\theta}(a_{t},G_{t}|s_{t},s_{G})=\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})p_{\theta}(G_{t}|s_{t},s_{G})$. We note that [17] found it helpful to also utilize a model-based return-conditioned policy, whereby the future state is modelled as part of the joint distribution being learned. This is shown to provide a useful regularizing signal for the policy, even though the future state prediction is not directly used at inference time. In this work, we also found it helpful to regularize the learned policy by predicting the full sequence of future states. The final distribution we are aiming to model is thus given by $p_{\theta}(s_{t+1:T},a_{t},G_{t}|s_{t},s_{G})=p_{\theta}(s_{t+1:T}|s_{t},s_{G},G_{t},a_{t})\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})p_{\theta}(G_{t}|s_{t},s_{G})$.
@@
At inference time, we obtain actions by first sampling returns $G_{t}\sim p_{\theta}(G_{t}|s_{t},s_{G})$ and then sampling actions $a_{t}\sim\pi_{\theta}(a_{t}|s_{t},s_{G},G_{t})$. This sampling procedure corresponds to the imitative policy since the sampled returns are obtained from the learned density that models the data distribution. Following prior work in offline RL [16, 17, 18], we can also sample actions from an exponentially-tilted policy distribution. This is done by sampling the returns from the tilted distribution $G^{\prime}_{t}\sim p_{\theta}(G_{t}|s_{t},s_{G})\exp(\kappa G_{t})$, with $G^{\prime}_{t}$ being the tilted return-to-go and where $\kappa$ represents the inverse temperature; higher values of $\kappa$ concentrate more density around the best outcomes or higher returns, while negative values of $\kappa$ concentrate on less favourable outcomes or lower returns.
@@
We are interested in modelling and controlling the individual components of the reward function rather than maximizing their weighted sum. For example, we would like to model an agent’s ability to reach its goal, drive on the road, and avoid collisions. In general, given $C$ reward components, our objective is to learn policies that are conditioned on all its factored dimensions as this would grant us control over each one at test time. This entails modelling separate return components as $G^{c}_{t}\sim p_{\theta}(G^{c}_{t}|s_{t},s_{G})$ for each return component $c$. Applying this factorization, we reformulate the learned policy to explicitly account for the conditioning on all return components $\pi_{\theta}(a_{t}|s_{t},s_{G},G^{1}_{t},\ldots G^{C}_{t})$. At test time, each return component will be accompanied by its own inverse temperature $\kappa^{c}$ to enable control over each return component, which enables sampling actions that adhere to different behaviours specified by $\{\kappa^{1},\ldots,\kappa^{C}\}$, as shown in Algorithm 2 in Appendix B.
@@
To implement our framework for behaviour simulation, we extend the approach presented above to the multi-agent setting. Across all agents we have sets for the joint states $\mathbb{S}_{t}$, goal states $\mathbb{S}_{G}$, actions $\mathbb{A}_{t}$, and returns-to-go $\mathbb{G}_{t}$. The final multi-agent joint distribution we model is:
@@ 2.2 Multi-Agent Behaviour Simulation Architecture
In this section, we introduce the proposed architecture for multi-agent behaviour simulation within the CtRL-Sim framework that parameterizes the multi-agent joint distribution presented in Equation (1). We propose an encoder-decoder Transformer architecture [19], as illustrated in Figure 2, where the encoder encodes the initial scene and the decoder autoregressively generates the trajectory rollout for all agents in the scene.
@@
Encoder To encode the initial scene, we first process the initial agent states and goals $(\mathbf{s}_{0},\mathbf{s}_{G})$ and the map context $m$, where $\mathbf{s}_{0}$ is the joint initial state of all agents and $\mathbf{s}_{G}$ is the joint goal state of all agents. Each agent $i$’s initial state information $s^{i}_{0}$, which includes the position, velocity, heading, and agent type, is encoded with an MLP. Similarly, each agent’s goal $s^{i}_{G}$, which is represented as the ground-truth final position, velocity, and heading, is also encoded with an MLP. We then concatenate the initial state and goal embedding of each agent and embed them with a linear layer to get per-agent embeddings of size $d$. We additionally apply an additive learnable embedding to encode the agents’ identities across the sequence of agent embeddings. The map context is encoded using a polyline map encoder, detailed more fully in Appendix F, which yields $L$ road segment embeddings of size $d$. The initial agent embeddings and road segment embeddings are then concatenated into a sequence of length $N+L$ and processed by a sequence of $E$ Transformer encoder blocks.
@@
Decoder The proposed decoder architecture models the joint distribution in Equation (1) as a sequence modelling problem, where we model the probability of the next token in the sequence conditioned on all previous tokens $p_{\theta}(x_{t}|x_{<t})$ [15]. In this work, we consider trajectory sequences of the form: $x=\langle\dots,(s^{1}_{t},s^{1}_{G}),(G^{1,1}_{t},\dots,G^{C,1}_{t}),a^{1}_{t},...,(s^{N}_{t},s^{N}_{G}),(G^{1,N}_{t},\dots,G^{C,N}_{t}),a^{N}_{t},\dots\rangle$. These sequences are an extension of the sequences considered in the Multi-Game Decision Transformer [16] to the multi-agent goal-conditioned setting with factorized returns. Unlike Decision Transformer [15], our model predicts the return distribution and samples from it at inference time, which enables flexible control over the agent behaviours and circumvents the need to specify an expert return-to-go. We obtain state-goal tuple $(s^{i}_{t},s^{i}_{G})$ embeddings in the same way that $(\mathbf{s}_{0},\mathbf{s}_{G})$ are processed in the encoder. Following recent work that tokenizes driving trajectories [20, 9], we discretize the actions and return-to-gos into uniformly quantized bins. We then embed the action and return-to-go tokens with a linear embedding. To each input token, we additionally add two learnable embeddings representing the agent identity and timestep, respectively. The tokenized sequence is then processed by $D$ Transformer decoder layers with a temporally causal mask that is modified to ensure that the model is permutation equivariant to the agent ordering (see Appendix F for details).
@@
Training Given a dataset of offline trajectories (Section 3), we train our model by sampling sequences of length $H\times N\times 3$, where $H$ is the number of timesteps in the context. The state, return-to-go, and action token embeddings output by the decoder are used to predict the next return token, action token, and future state sequence, respectively. We train the return-to-go and action headers with the standard cross-entropy loss function and the future state sequence header with an L2 regression loss function. The final loss function is of the form: $\mathcal{L}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{return-to-go}}+\alpha\mathcal{L}_{\text{state}}$.